Caitlyn got a 559-peso plan for her new cellphone. This plan covers all calls and texts to Globe, but none for other networks. It costs a peso per minute per text to other networks. Write the equation that signifies this relationship. Let Y be the amount of money to be paid per month and X the number of texts to other networks. How much does she have to pay if she sent 195 texts to Sun and 120 texts to Smart?
Answer by Theo(13342)

she pays 559 a month plus 1 peso per text to other networks besides globe.

formula is y = 559 + 1 * x

x is the number of texts to networks other than globe.
the charge is 1 peso per text.

if she sent 195 texts to sun and 120 texts to smart, then she had a total of 315 texts to networks other than globe.

her monthly charges would be y = 559 + 1 * (120 + 195) = 559 + 1 * (315) = 559 + 315 = 874.